1.7. Problem. (Sections 2.2-2.4)

(a) A quiz requires the student to match 10 terms with their correct definitions, where 10 definitions are given. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.

(b) A quiz requires the student to match 10 terms with their correct definitions, where 11 definitions are given and one of the definitions does not match any of the terms. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.

(c) A quiz requires the student to match 10 terms with their correct definitions, where 12 definitions are given and two of the definitions do not match any of the terms. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.
Answer #1

(a) A quiz requires the student to match 10 terms with their correct definitions, where 10 definitions are given. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.

Probability of success, p = 1/10 = 0.1

n = 10

X P(X)
0 0.34868
1 0.38742
2 0.19371
3 0.05740
4 0.01116
5 0.00149
6 0.00014
7 0.00001
8 0.00000
9 0.00000
10 0.00000

Required probability = 0.00001 + 0.00000 + 0.00000 + 0.00000 = 0.00001

(b) A quiz requires the student to match 10 terms with their correct definitions, where 11 definitions are given and one of the definitions does not match any of the terms. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.

Probability of success, p = 1/11 = 0.090909

n = 10

X P(X)
0 0.38554
1 0.38554
2 0.17349
3 0.04627
4 0.00810
5 0.00097
6 0.00008
7 0.00000
8 0.00000
9 0.00000
10 0.00000

Required probability = 0.00000 + 0.00000 + 0.00000 + 0.00000 = 0.00000

(c) A quiz requires the student to match 10 terms with their correct definitions, where 12 definitions are given and two of the definitions do not match any of the terms. If a student is just guessing, find the probability that they match at least 7 terms with their correct definitions.

Probability of success, p = 1/12 = 0.0833

n = 10

X P(X)
0 0.41890
1 0.38082
2 0.15579
3 0.03777
4 0.00601
5 0.00066
6 0.00005
7 0.00000
8 0.00000
9 0.00000
10 0.00000

Required probability = 0.00000 + 0.00000 + 0.00000 + 0.00000 = 0.00000